# MotionLab: Unified Human Motion Generation and Editing via the Motion-Condition-Motion Paradigm

### v4

@@ Abstract.
Human motion generation and editing are key components of computer vision. However, current approaches in this field tend to offer isolated solutions tailored to specific tasks, which can be inefficient and impractical for real-world applications. While some efforts have aimed to unify motion-related tasks, these methods simply use different modalities as conditions to guide motion generation. Consequently, they lack editing capabilities, fine-grained control, and fail to facilitate knowledge sharing across tasks. To address these limitations and provide a versatile, unified framework capable of handling both human motion generation and editing, we introduce a novel paradigm: Motion-Condition-Motion, which enables the unified formulation of diverse tasks with three concepts: source motion, condition, and target motion. Based on this paradigm, we propose a unified framework, MotionLab, which incorporates rectified flows to learn the mapping from source motion to target motion, guided by the specified conditions. In MotionLab, we introduce the 1) MotionFlow Transformer to enhance conditional generation and editing without task-specific modules; 2) Aligned Rotational Position Encoding to guarantee the time synchronization between source motion and target motion; 3) Task Specified Instruction Modulation; and 4) Motion Curriculum Learning for effective multi-task learning and knowledge sharing across tasks. Notably, our MotionLab demonstrates promising generalization capabilities and inference efficiency across multiple benchmarks for human motion. Our code and additional video results are available at: Project Website.
Motion Planning, Motion Control
journal: TOGccs: Computing methodologies Motion processingccs: Computing methodologies Motion captureccs: Computing methodologies Virtual reality
@@
Figure 1. Demonstration of our MotionLab’s versatility, performance and efficiency. Previous SOTA refer to multiple expert models, including MotionLCM (Dai et al., 2025), OmniControl (Xie et al., 2023), MotionFix (Athanasiou et al., 2024), CondMDI (Cohan et al., 2024) and MCM-LDM (Song et al., 2024). All motions are represented using SMPL (Loper et al., 2023), where transparent motion indicates the source motion or condition, and the other represents the target motion. More qualitative results are available in the website and appendix.
1. Introduction
@@
Human motion is a crucial component of computer vision, with applications spanning game development, film production, and virtual reality (Guo et al., 2022a; Tevet et al., 2023). With the advancements of generative diffusion models (Ho et al., 2020; Song et al., 2020; Dhariwal and Nichol, 2021), human motion generation has garnered considerable attention, aiming at generating human motion aligned with the input conditions, such as text (Tevet et al., 2023; Zhang et al., 2022; Tevet et al., 2022) and trajectory (i.e., joints’ coordinates) (Xie et al., 2023; Fujiwara et al., 2025; Sun et al., 2024; Dai et al., 2025; Athanasiou et al., 2023, 2022; Zhang et al., 2023a; Song et al., 2023; Sampieri et al., 2024). On the other hand, in order to modify the motion assets in the industry, significant efforts have been dedicated to motion editing, which intended to modify the properties of prepared motion like motion style transfer (Zhong et al., 2025; Song et al., 2024; Aberman et al., 2020; Guo et al., 2024b; Jang et al., 2022).
As summarized in Table 1, current research in this domain mainly develops isolated, task-specific solutions, forcing practitioners to train multiple models for human motion generation and editing—a strategy that is inefficient and impractical. Although several studies (Shrestha et al., 2025; Zhou and Wang, 2023; Zhou et al., 2023; Zhang et al., 2025a; Yang et al., 2024; Luo et al., 2024; Fan et al., 2024) have attempted to unify motion-related tasks, they merely consider different modalities as generation conditions, leading to limited editing capabilities and insufficient fine-grained trajectory control. Moreover, these approaches overlook the intrinsic links between motion generation and editing, thereby hindering potential knowledge sharing. In contrast, a well-designed unified framework can exploit the large volumes of multi-task data to potentially surpass specialist models through effective cross-task representation learning. Motivated by this prospect and inspired by the success of large language models in unifying diverse tasks (Achiam et al., 2023; Dubey et al., 2024), we pose the following question: Can human motion generation and editing be effectively unified within a single framework?
In response to this question, we aim to design an elegant and scalable paradigm. Drawing inspiration from the next-token prediction paradigm (Brown et al., 2020; Achiam et al., 2023), which has revolutionized the NLP field, we propose a novel paradigm: Motion-Condition-Motion. This paradigm is built upon three concepts – source motion, condition, and target motion. Similar to how next-token prediction anticipates the subsequent word based on context, the Motion-Condition-Motion paradigm predicts the target motion based on the source motion and specified conditions. For any human motion generation task, the source motion can be treated as none, and the target motion must align with the provided conditions. For any human motion editing task, the target motion is derived from the source motion based on the conditions. By unifying these tasks within this elegant and scalable paradigm, this framework can be seamlessly extended to various human motion tasks and scaled across diverse datasets. Given that human motions are inherently tied to their semantics, trajectories, and styles in practical applications, we aim to unify several key tasks under this framework. These tasks include text-based motion generation and editing (Tevet et al., 2023; Zhang et al., 2022; Tevet et al., 2022; Athanasiou et al., 2024; Goel et al., 2024), trajectory-based motion generation and editing (Xie et al., 2023; Dai et al., 2025), motion in-between (Cohan et al., 2024; Harvey et al., 2020) and motion style transfer (Song et al., 2024; Zhong et al., 2025), as illustrated in Figure 1.
@@
MDM (Tevet et al., 2023) $\checkmark$ $\times$ $\times$ $\times$ -- $\times$
MLD (Chen et al., 2023) $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\times$
OmniControl (Xie et al., 2023) $\checkmark$ $\times$ $\checkmark$ $\times$ -- $\times$
MotionFix (Athanasiou et al., 2024) -- $\checkmark$ $\times$ $\times$ -- $\times$
CondMDI (Cohan et al., 2024) $\checkmark$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\times$
MCM-LDM (Song et al., 2024) $\times$ $\times$ $\times$ $\times$ -- $\checkmark$
MotionGPT (Jiang et al., 2023) $\checkmark$ -- $\times$ $\times$ $\checkmark$ $\times$
MotionCLR (Chen et al., 2024) $\checkmark$ -- $\times$ $\times$ -- --
@@
Table 1. Summary of different methods focusing on motion generation and editing. $\checkmark$ indicates that the method has been trained for the task, $\times$ indicates that the method has not been trained, and -- indicates that the method has not been trained but can be implemented in a zero-shot manner.
@@
Despite the proposed paradigm, several significant challenges remain in balancing versatility, performance, and efficiency: 1) Unifying various tasks inevitably introduces additional modalities, while each modality may involve multiple tasks. A naive solution, like adopting multiple cross-attention mechanisms for each task in generation-unified frameworks (Fan et al., 2024; Zhang et al., 2025a), is suboptimal. 2) More sampling time is required for certain tasks (e.g., trajectory-based motion generation and motion in-between (Xie et al., 2023; Zhong et al., 2025)), as existing methods in these areas involve task-specific posterior guidance (Chung et al., 2022) during inference to improve conditional guidance. 3) Time asynchrony between the source motion and target motion may arise due to the limited scale of the paired editing dataset and the use of implicit positional encoding (Chen et al., 2023; Xie et al., 2023; Athanasiou et al., 2024; Cohan et al., 2024). 4) Most importantly, naively integrating various motion generation and editing tasks into a single framework could lead to task conflicts and catastrophic forgetting, impairing the framework’s overall performance.
@@
To address these challenges, we propose a novel generative framework, named MotionLab, built upon our designed MotionFlow Transformer (MFT) as shown in Figure 3. Inspired by MM-DiT (Esser et al., 2024), our MFT also leverages rectified flows (Liu et al., 2022; Lipman et al., 2022), but we utilize them to map source motion to target motion based on specified conditions. Unlike MM-DiT that focuses exclusively on text and images, our MFT incorporates multiple modalities: source motion, target motion, text, trajectory, and style. In MFT, each modality is allocated a dedicated modality path and fully interacts with the others through joint attention, which enables MFT to enhance conditional generation and editing without requiring task-specific modules or posterior guidance for certain tasks. To ensure temporal synchronization between source motion and target motion, we incorporate an Aligned ROtational Position Encoding (Aligned ROPE) into MFT, explicitly aligning tokens in corresponding frames between the source and target motion. Additionally, to adapt one modality to different tasks, we propose Task Instruction Modulation, which distinguishes different tasks for each modality by introducing an additional task embedding into the MFT. To harmoniously integrate diverse tasks, we propose a curriculum-inspired training strategy termed Motion Curriculum Learning based on the easy-to-hard training principle. Intuitively, tasks involving fewer modalities or more explicit conditional information (e.g., source motion) present lower complexity and are therefore prioritized in the learning sequence.
@@
In this paper, motion-related tasks are decomposed into combinations of modalities through the Motion-Condition-Motion paradigm. These modalities are subsequently represented via each modality paths within the MFT, learning cross-modal interactions through the joint attention, while adapting individual modalities to different tasks through Task Instruction Modulation. By implementing a curriculum learning from single to multiple, from simple (e.g., source motion and trajectory) to complex (e.g., text and style) modalities, the spatial knowledge inherent in 3D representations can be effectively transferred to more abstract modalities since the modalities of the former can represent the latter. Through these designs, we validate MotionLab on multiple benchmarks, demonstrating superior versatility, performance, and efficiency compared to baselines across various human motion generation and editing tasks.
2. Related Works
@@
Motion Generation and Editing. Motion generation can be classified based on input conditions. Among these, text-based motion generation is one of the most compelling areas (Zhang et al., 2022; Tevet et al., 2022, 2023; Guo et al., 2022a; Chen et al., 2023; Guo et al., 2024a, 2022b; Wang et al., 2024b; Jiang et al., 2023; Kim et al., 2023; Lin et al., 2023; Lu et al., 2023; Plappert et al., 2016; Zhang et al., 2023b; Guo et al., 2020; Petrovich et al., 2021), as it trains models to comprehend the semantics of text and generate corresponding pose sequences. To address the fine-grained requirements of practical applications, trajectory-based motion generation has been proposed (Karunratanakul et al., 2023; Shafir et al., 2023; Xie et al., 2023; Dai et al., 2025; Zhang et al., 2023a), where specific motion properties, such as joints reaching designated positions at specified times, are defined. Additionally, motion in-between (Tevet et al., 2023; Jiang et al., 2023; Cohan et al., 2024; Qin et al., 2022; Pinyoanuntapong et al., 2024) focuses on generating complete motion sequences given key poses at keyframes. To enable in-place editing of human motion (Goel et al., 2024; Athanasiou et al., 2024), MotionFix (Athanasiou et al., 2024) introduces text-based motion editing using paired source and target motions. We extend this approach to trajectory-based motion editing by substituting text with joint trajectories. Meanwhile, style plays a crucial role in human motion, leading to motion style-transfer (Jang et al., 2022; Aberman et al., 2020; Zhong et al., 2025; Song et al., 2024). However, the aforementioned methods concentrate solely on specific tasks, rendering them impractical for real-world applications. Moreover, they overlook the intrinsic connections across different human motion tasks and fail to facilitate knowledge sharing among these tasks. In contrast, our unified framework enhances performance on data-scarce editing tasks through multi-task learning.
@@
Unified frameworks for human motion. There are also some efforts in existing methods that try to unify tasks related to human motion. One line of work (Jiang et al., 2023, 2025; Zhou et al., 2024; Li et al., 2024; Wang et al., 2024b; Ling et al., 2024; Wu et al., 2024; Luo et al., 2024; Athanasiou et al., 2024) focuses on motion understanding, such as motion captioning or describing human motion in images and videos. Yet, these approaches often rely on GPT-like structures, which requires a large amount of training resources and GPU memory. In addition, they fail to provide fine-grained control (e.g., trajectory-based generation and editing) over motion, which is crucial in practical applications. Another line of effort (Shrestha et al., 2025; Zhou and Wang, 2023; Zhou et al., 2023; Zhang et al., 2025a; Yang et al., 2024; Luo et al., 2024; Fan et al., 2024; Alexanderson et al., 2023) highlights generating motion based on more modalities, such as music and speech. However, these approaches just only integrate more modalities into one model and cannot flexibly edit motion, which can cause them suffering from the multi-task learning and limit their scope of use. The closest to our work are FLAME (Kim et al., 2023) and MotionCLR (Chen et al., 2024). However, FLAME does not support style transfer and precise text-based editing like “move faster”, and MotionCLR does not support trajectory-based generation and editing, requiring cumbersome manual adjustments to the attention.
3. PRELIMINARY: Rectified Flows
Figure 2. Demonstration of the difference trajectory between diffusion models and rectified flows. This difference lies in that the trajectory of diffusion models is based on $x_{t}=\sqrt{(1-\overline{\alpha_{t}})}x_{0}+\sqrt{\overline{\alpha_{t}}}\epsilon$, while the trajectory of rectified flows is based on $x_{t}=(1-t)x_{0}+tx_{1}$. This distinction leads to more robust learning by maintaining a constant velocity, contributing to model’s efficiency (Zhao et al., 2024).
@@ 3. PRELIMINARY: Rectified Flows
Flow-based methods (Liu et al., 2022; Lipman et al., 2022; Esser et al., 2024; Ma et al., 2024; Fei et al., 2024; Polyak et al., 2024) have recently received significant attention due to their generalizability and efficiency compared to diffusion models. Specifically, these methods directly regress the transport vector field between noise distribution $p_{1}$ and data distribution $p_{0}$ with the straightest possible trajectories and sample by the corresponding ordinary differential equation (ODE) (Wang et al., 2024a). Among these methods, rectified flows (Liu et al., 2022; Lipman et al., 2022) aim to learn a trajectory from noise $x_{0}$ to data $x_{1}$, which can be formulated as $x_{t}=\varphi(x_{0},x_{1},t)$, and the velocity field $v_{t}$ of the trajectory $x_{t}$ can be defined by:
(1) $v_{t}=\frac{dx_{t}}{dt}=\frac{\partial\varphi_{t}(x_{0},x_{1},t)}{\partial t},t\in[0,1]$
@@
Once we have learned this velocity field $v_{t}$, we can get $x_{0}$ from any $x_{1}$ by numerically integrating. Hence, rectified flows $v_{\theta}$ are trained to predict $v_{t}$ by given $x_{t}$ and $t$, and the training object of rectified flows can be represented as:
@@
(3) $x_{t}=(1-t)x_{0}+tx_{1}$
(4) $v_{t}=x_{1}-x_{0}$
Therefore, the training objective can be reformulated as:
[…]
(6) $x_{t-\frac{1}{N}}=x_{t}+\frac{1}{N}v_{\theta}(t,x_{t})$
where $N$ is the discretization number of the interval [0,1].
4. Motion-Condition-Motion
To unify the tasks of human motion generation and editing in an elegant and scalable paradigm, we propose the paradigm of Motion-Condition-Motion. As shown in Table 2, all these tasks are unified by three concepts: source motion, condition, and target motion.

### v5

@@ Abstract
Human motion generation and editing are key components of computer vision. However, current approaches in this field tend to offer isolated solutions tailored to specific tasks, which can be inefficient and impractical for real-world applications. While some efforts have aimed to unify motion-related tasks, these methods simply use different modalities as conditions to guide motion generation. Consequently, they lack editing capabilities, fine-grained control, and fail to facilitate knowledge sharing across tasks. To address these limitations and provide a versatile, unified framework capable of handling both human motion generation and editing, we introduce a novel paradigm: Motion-Condition-Motion, which enables the unified formulation of diverse tasks with three concepts: source motion, condition, and target motion. Based on this paradigm, we propose a unified framework, MotionLab, which incorporates rectified flows to learn the mapping from source motion to target motion, guided by the specified conditions. In MotionLab, we introduce the 1) MotionFlow Transformer to enhance conditional generation and editing without task-specific modules; 2) Aligned Rotational Position Encoding to guarantee the time synchronization between source motion and target motion; 3) Task Specified Instruction Modulation; and 4) Motion Curriculum Learning for effective multi-task learning and knowledge sharing across tasks. Notably, our MotionLab demonstrates promising generalization capabilities and inference efficiency across multiple benchmarks for human motion. Our code and additional video results are available at: https://diouo.github.io/motionlab.github.io/.
@@
Figure 1: Demonstration of our MotionLab’s versatility, performance and efficiency. Ours specialists refer to the proposed framework tailored for specified tasks. Previous SOTA refer to multiple models, including MotionLCM [12], OmniControl [64], MotionFix [6], CondMDI [11] and MCM-LDM [55]. All motions are represented using SMPL [39], where transparent motion indicates the source motion or condition, and the other represents the target motion. More qualitative results are available in the website and appendix.
1 Introduction
@@
Human motion is a crucial component of computer vision, with applications spanning game development, film production, and virtual reality [21, 59]. With the advancements of generative diffusion models [28, 53, 13], human motion generation has garnered considerable attention, aiming at generating human motion aligned with the input conditions, such as text [59, 67, 58] and trajectory (i.e., joints’ coordinates) [64, 18, 57, 12, 5, 4, 68, 54, 50]. Concurrently, to maximize the utility of motion assets within industry settings, significant efforts have been dedicated to motion editing tasks, including motion style transfer [71, 55, 1, 24, 29].
@@
MDM [59] $\checkmark$ $\times$ $\times$ $\times$ -- $\times$
MLD [9] $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\times$
OmniControl [64] $\checkmark$ $\times$ $\checkmark$ $\times$ -- $\times$
MotionFix [6] -- $\checkmark$ $\times$ $\times$ -- $\times$
CondMDI [11] $\checkmark$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\times$
MCM-LDM [55] $\times$ $\times$ $\times$ $\times$ -- $\checkmark$
MotionGPT [30] $\checkmark$ -- $\times$ $\times$ $\checkmark$ $\times$
MotionCLR [8] $\checkmark$ -- $\times$ $\times$ -- --
@@
Table 1: Summary of methods focusing on motion generation and editing. $\checkmark$ indicates that the method has been trained for the task, $\times$ indicates that the method fails to implement, and -- indicates that the method has not been trained but can implement in a zero-shot manner.
As summarized in Table 1, current research in this domain mainly develops task-specific solutions, forcing practitioners to train multiple models for human motion generation and editing, which is inefficient and impractical. Although several studies [52, 72, 73, 69, 65, 41, 16] have attempted to unify motion-related tasks, they merely consider different modalities as generation conditions, leading to limited editing capabilities and insufficient fine-grained trajectory control. Moreover, these approaches overlook the intrinsic links between motion generation and editing, thereby hindering potential knowledge sharing. In contrast, a well-designed unified framework can exploit the large volumes of multi-task data to surpass specialist models through effective cross-task representation learning. Motivated by this prospect and inspired by the success of large language models in unifying NLP tasks [2, 14], we pose the following question: Can human motion generation and editing be effectively unified within a single framework?
In response to this question, it is essential to design an elegant and scalable paradigm. Hence, we propose a novel paradigm: Motion-Condition-Motion. This paradigm is built upon three concepts – source motion, condition, and target motion. Concretely, the target motion is predicted by the source motion and specified conditions in this Motion-Condition-Motion paradigm. For any human motion generation task, the source motion can be treated as none, and the target motion must align with the provided conditions. For any human motion editing task, the target motion is derived from the source motion based on the conditions. By unifying these tasks within this elegant and scalable paradigm, this framework can be seamlessly extended to various human motion tasks and scaled across diverse datasets. Given that human motions are inherently tied to their semantics, trajectories, and styles in practical applications, we aim to unify several key tasks under this framework. These tasks include text-based motion generation and editing [59, 67, 58, 6, 19], trajectory-based motion generation and editing [64, 12, 25], motion in-between [11, 26] and motion style transfer [55, 71], as illustrated in Figure 1.
@@
Despite the proposed paradigm, several significant challenges remain in balancing versatility, performance, and efficiency: 1) Unifying various tasks inevitably introduces additional modalities, while each modality may involve multiple tasks. A naive solution, like adopting multiple cross-attention mechanisms for each task in generation-unified frameworks [16, 69], is suboptimal. 2) More sampling time is required for certain tasks (e.g., trajectory-based motion generation and motion in-between [64, 71]), as existing methods in these areas involve task-specific posterior guidance [10] during inference to improve conditional guidance. 3) Time asynchrony between the source motion and target motion may arise due to the limited scale of the paired editing dataset and the use of implicit positional encoding [9, 64, 6, 11]. 4) Most importantly, naively integrating various motion generation and editing tasks into a single framework could lead to task conflicts and catastrophic forgetting, impairing the framework’s overall performance.
@@
To address these challenges, we propose a novel generative framework, termed MotionLab, built upon rectified flows [38, 37] and MM-DiT [15], as illustrated in Figure 2. Rectified flows are particularly well-suited for MotionLab since they are designed to implement the optimal transport between source and target distributions, naturally aligning with the Motion-Condition-Motion paradigm. Furthermore, human motion must adhere to skeletal kinematics. Consequently, the distribution of valid target motions is highly restricted, and individual target motions may correspond to multiple conditions (i.e., many-to-one), which suggests that a shared transport map can be efficiently transferred across tasks. In contrast to MM-DiT, our proposed Motion Flow Transformer (MFT) encompasses more modalities, including source motion, target motion, text, trajectory, and style. Within the MFT, each modality is assigned a dedicated path, and comprehensive interaction among modalities is facilitated via joint attention. This architecture enables MFT to advance conditional generation and editing capabilities without necessitating task-specific modules or posterior guidance for particular tasks. To ensure temporal synchronization between source and target motions, we incorporate an Aligned Rotational Position Encoding into MFT, which explicitly aligns tokens at corresponding frames between the source and target sequences. Moreover, to enable adaptation of a single modality to various tasks, we introduce Task Instruction Modulation, which flexibly embeds various tasks into the MFT. To seamlessly integrate diverse tasks, we propose a curriculum-inspired training strategy, termed Motion Curriculum Learning, based on an easy-to-hard training principle.
@@
In this paper, motion generation and editing are decomposed into combinations of modalities through the Motion-Condition-Motion paradigm. These modalities are subsequently represented via each modality’s paths within the MFT, learning inter-modal interactions through the joint attention, while adapting individual modalities to different tasks through Task Instruction Modulation. By implementing a curriculum learning from single to multiple, from simple (e.g., source motion and trajectory) to complex (e.g., text and style) modalities, the spatial knowledge inherent in 3D representations can be effectively transferred to the latter modalities since the modalities of the former can represent the latter. Through these designs, we validate MotionLab on multiple benchmarks, demonstrating superior versatility, performance, and efficiency compared to baselines across various human motion generation and editing tasks.
2 Related Work
@@
Motion Generation and Editing. Motion generation can be classified based on input conditions. Among these, text-based motion generation is one of the most compelling areas [67, 58, 59, 21, 9, 23, 22, 62, 30, 33, 35, 40, 46, 66, 20, 44], as it trains models to comprehend the semantics of text and generate corresponding pose sequences. To address the fine-grained requirements of practical applications, trajectory-based motion generation has been proposed [32, 51, 64, 12, 68], where specific motion properties, such as joints reaching designated positions at specified times, are defined. Additionally, motion in-between [59, 30, 11, 48, 45] focuses on generating complete motion sequences given key poses at keyframes. To enable in-place editing of human motion [19, 6], MotionFix [6] introduces text-based motion editing using paired source and target motions. We extend this approach to trajectory-based motion editing by substituting text with joint trajectories. Meanwhile, style plays a crucial role in human motion, leading to motion style-transfer [29, 1, 71, 55]. However, the aforementioned methods concentrate solely on specific tasks, rendering them impractical for real-world applications. Moreover, they overlook the intrinsic connections across different human motion tasks and fail to facilitate knowledge sharing among these tasks. In contrast, our unified framework enhances performance on data-scarce editing tasks through multi-task learning.
@@
Unified Frameworks for Human Motion. There are also some efforts in existing methods that try to unify tasks related to human motion. One line of work [30, 31, 74, 34, 62, 36, 63, 41, 6] focuses on motion understanding, such as motion captioning or describing human motion in images and videos. Yet, these approaches often rely on GPT-like structures, which require a large amount of training resources and GPU memory. In addition, they fail to provide fine-grained control (e.g., trajectory-based generation and editing) over motion, which is crucial in practical applications. Another line of effort [52, 72, 73, 69, 65, 41, 16, 3] highlights generating motion based on more modalities, such as music and speech. However, these approaches only integrate more modalities into one model and cannot flexibly edit motion, which can cause them to suffer from multi-task learning and limit their scope of use. The closest to our work are FLAME [33] and MotionCLR [8]. However, FLAME does not support style transfer and precise text-based editing like “move faster”, and MotionCLR does not support trajectory-based generation and editing, requiring cumbersome manual adjustments to the attention.
3 Preliminary: Rectified Flows
@@
Flow-based generative methods [38, 37, 15, 42, 17, 47] have recently received significant attention due to their generalizability and efficiency compared to diffusion models. Specifically, these methods directly regress the transport vector field between the source distribution $p_{1}$ and target distribution $p_{0}$ with the straightest possible trajectories and sample by the corresponding ordinary differential equation (ODE) [61]. Among these methods, rectified flows [38, 37] aim to learn a trajectory from source data $x_{0}$ to target data $x_{1}$, which can be formulated as $x_{t}=\varphi(x_{0},x_{1},t)$, and the velocity field $v_{t}$ of the trajectory $x_{t}$ can be defined by:
$v_{t}=\frac{dx_{t}}{dt}=\frac{\partial\varphi_{t}(x_{0},x_{1},t)}{\partial t},t\in[0,1]$ (1)
@@
Once we have learned this velocity field $v_{t}$, we can get $x_{0}$ from any $x_{1}$ by numerically integrating:
$x_{t-\frac{1}{N}}=x_{t}-\frac{1}{N}v_{\theta}(t,x_{t})$ (2)
where $N$ is the discretization number of the interval [0,1]. Hence, rectified flows $v_{\theta}$ are trained to predict $v_{t}$ by given $x_{t}$ and $t$, and the training objective can be represented as:
@@
4 Motion-Condition-Motion
To unify the tasks of human motion generation and editing, we propose the paradigm of Motion-Condition-Motion. As shown in Table 2, all these tasks are unified by three concepts: source motion, condition, and target motion.